### v1

@@ 5 Conclusions
We present FAST observations of the OH 18-cm lines in comet C/2025 A6 obtained between 23rd October and 8th November 2025. Our main results and conclusions can be summarized as follows:
1. We confirm that the detected OH 18-cm signals originate from the cometary coma. This conclusion is supported by (i) the agreement between the fitted line-center velocities and radial velocity of the comet relative to the telescope, and (ii) the On–Off observations, in which no emission is detected at the offset positions along the comet’s trajectory. We detected the OH main lines in emission on 23rd and 26th October and in absorption on 3rd to 5th November. The observed emission-to-absorption transition is consistent with the dependence of the OH inversion parameter on heliocentric radial velocity through the Swings effect, which can drive the inversion parameter across zero.
@@ Acknowledgements.
This work is supported by the National Key R & D Program of China No. 2025SKA0140100, No. 2022YFC2205202, No. 2020SKA0120100, and the National Natural Science Foundation of China (NSFC, Grant Nos. 12373032, 12003047, 11773041, U2031119, 12173052, and 12173053). Both Lei Qian and Zhichen Pan were supported by the Youth Innovation Promotion Association of CAS (id. 2018075, Y2022027, and 2023064) and the CAS “Light of West China” Program. Hongfei Liu has been supported by the National Natural Science Foundation of China (NSFC) under No.12273072. Liyun Zhang has been supported by the Science and Technology Program of Guizhou Province under project No.QKHPTRC-ZDSYS[2023]003 and QKHFQ[2023]003. FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences. This work made use of the data from FAST (Five-hundred-meter Aperture Spherical radio Telescope). We thank Prof. Oleg Smirnov and his group at MeerKAT for very helpful discussions. We also acknowledge the support of the University of Chinese Academy of Sciences and the use of its 70 cm telescope.Appendix A From Antenna temperature to Flux for FAST UWB data
@@
Here, represents the telescope’s efficiency, and is the theoretical gain based on the geometrical illumination area. We adopt the value of the Gain at the 1650 MHz from (2023RAA....23g5016Z), according to the zenith angle (ZA) which they detected derive the corresponding 19.6 K Jy-1, then get the actual gain through the relationship between ZA and (2019SCPMA..6259502J). The result was shown in Table 2.
@@ Appendix B The OH production rate with quenching
When considering the quenching effect, we assume that the inversion factor becomes for , so that OH radicals within this region do not contribute to the observable 18-cm signal (1981A&A....99..320D), and the factor can be written as:
Where the is the column density distribution after subtracted the sphere with radius . The quenching radius can be expressed as = (e.g., 2025A&A...701A.204L). Adopting following 2025A&A...701A.204L results in unstable convergence. We therefore adopt a reduced value of to ensure convergence. With this treatment, the derived values for 23rd October, 26th October, 4th November, and 5th November are , , , and s-1, respectively.

### v2

@@ 5 Conclusions
We present FAST observations of the OH 18-cm lines in comet C/2025 A6 obtained between October 23 and November 8. Our main results and conclusions can be summarized as follows:
1. We confirm that the detected OH 18-cm signals originate from the cometary coma. This conclusion is supported by (i) the agreement between the fitted line-center velocities and topocentric radial velocity of the comet, and (ii) the On–Off observations, in which no emission is detected at the offset positions along the comet’s trajectory. We detected the OH main lines in emission on October 23, 26 and in absorption on November 3 to 5. The observed emission-to-absorption transition is consistent with the dependence of the OH inversion parameter on heliocentric radial velocity through the Swings effect, which can drive the inversion parameter across zero.
@@ Acknowledgements.
We would like to thank the anonymous reviewers for their constructive suggestions, which have greatly improved the quality of this paper. This work is supported by the National Key R & D Program of China No. 2025SKA0140100, No. 2022YFC2205202, No. 2020SKA0120100, No. 2025SKA0140101 and the National Natural Science Foundation of China (NSFC, Grant Nos. 11703047, 12373032, 12003047, 11773041, U2031119, 12173052, and 12173053). Both Lei Qian and Zhichen Pan were supported by the Youth Innovation Promotion Association of CAS (id. 2018075, Y2022027, and 2023064) and the CAS “Light of West China” Program. Hongfei Liu has been supported by the National Natural Science Foundation of China (NSFC) under No.12273072. Liyun Zhang has been supported by the Science and Technology Program of Guizhou Province under project No. QKHPTRC-ZDSYS[2023]003 and QKHFQ[2023]003, as well as the Guizhou Provincial Natural Science Foundation project NO. ZD[2026]058. FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences. This work made use of the data from FAST. We thank Prof. Oleg Smirnov and his group at MeerKAT for very helpful discussions. We also acknowledge the support of the University of Chinese Academy of Sciences and the use of its 70 cm telescope.References
- Comets: looking ahead. Philosophical Transactions of the Royal Society of London Series A 375 (2097), pp. 20160261. External Links: Document Cited by: §1.
- A Study of the Comets with Large Perihelion Distances C/2019 L3 (ATLAS) and C/2019 O3 (Palomar). Research in Astronomy and Astrophysics 24 (9), pp. 095018. External Links: Document Cited by: §1.
- OH Observation of Comet Kohoutek (1973f) at 18 CM Wavelength. A&A 34, pp. 163. Cited by: §1.
- Retrieving the coma gas expansion velocity in P/Halley, Wilson (1987 VII) and several other comets from the 18-cm OH line shapes. A&A 238 (1-2), pp. 382–400. Cited by: §1, §4.2, §4.3, §4.3.
- The composition of cometary volatiles. In Comets II, M. C. Festou, H. U. Keller, and H. A. Weaver (Eds.), pp. 391. Cited by: §1.
- Radio observations of the hydroxyl radical in comets with high spectral resolution - Kinematics and asymmetries of the OH coma in C/Meier (1978 XXI), C/Bradfield (1979X), and C/Austin (1982g). A&A 131 (1), pp. 111–122. Cited by: §4.2.
- FAST Observations of Four Comets to Search for the Molecular Line Emissions between 1.0 and 1.5 GHz Frequencies. Research in Astronomy and Astrophysics 24 (10), pp. 105008. External Links: Document, 2409.06227 Cited by: §1.
- Observations at Nançay of the OH 18-cm lines in comets. The data base. Observations made from 1982 to 1999. A&A 393, pp. 1053–1064. External Links: Document Cited by: §1, §4.2, §4.3.
- The photodissociation of water in cometary atmospheres. A&A 213 (1-2), pp. 459–464. Cited by: §1, §4.2.
- The OH radiacal in comets: observation and analysis of the hyperfine microwave transitions at 1667 MHz and 1665 MHz.. A&A 99, pp. 320–340. Cited by: Appendix B, §1, §4.1.
- Low NH3/H2O ratio in comet C/2020 F3 (NEOWISE) at 0.7 au from the Sun. A&A 677, pp. A157. External Links: Document, 2307.11486 Cited by: §4.2.
- Comet C/2025 a6 (lemmon). Central Bureau Electronic Telegrams 5508, pp. 1. Cited by: §1.
- Solar Photo Rates for Planetary Atmospheres and Atmospheric Pollutants. Ap&SS 195 (1), pp. 1–294. External Links: Document Cited by: §4.3.
- Commissioning progress of the FAST. Science China Physics, Mechanics, and Astronomy 62 (5), pp. 959502. External Links: Document, 1903.06324 Cited by: Appendix A.
- Comet Taxonomy. In Completing the Inventory of the Solar System, T. Rettig and J. M. Hahn (Eds.), Astronomical Society of the Pacific Conference Series, Vol. 107, pp. 173–191. Cited by: §1.
- Pre-perihelion radio observations of comet 12P/Pons-Brooks with the Tianma Radio Telescope. A&A 701, pp. A204. External Links: Document, 2508.06936 Cited by: Appendix B, §4.2, §4.3.
- ALMA and GMRT Studies of Dust Continuum Emission and Spectral Lines Toward Oort Cloud Comet C/2022 E3 (ZTF). Research in Astronomy and Astrophysics 24 (12), pp. 125009. External Links: Document, 2411.03049 Cited by: §1.
- The Five-Hundred Aperture Spherical Radio Telescope (fast) Project. International Journal of Modern Physics D 20 (6), pp. 989–1024. External Links: Document, 1105.3794 Cited by: §2.
- Comets in the Post-Halley ERA - Volume 1; Volume 2. Astrophysics and Space Science Library, Vol. 167. External Links: Document Cited by: §1.
- A radio continuum survey of the northern sky at 1420 MHz. II. A&AS 63, pp. 205. Cited by: §4.3.
- A radio continuum survey of the northen sky at 1420 MHz - Part I.. A&AS 48, pp. 219–297. Cited by: §4.3.
[…]
- OH fluorescence in comets - Fluorescence efficiency of the ultraviolet bands. ApJ 258, pp. 864–877. External Links: Document Cited by: §4.1.
- The Fluorescence of Cometary OH. ApJ 331, pp. 1058. External Links: Document Cited by: 1st item, §4.1, §4.3.
- Models of cometary emission in the 18-cm OH transitions - The predicted behavior of Comet Halley. AJ 90, pp. 1117–1135. External Links: Document Cited by: §4.3.
- A radio continuum survey of the southern sky at 1420 MHz. Observations and data reduction. A&A 368, pp. 1123–1132. External Links: Document, astro-ph/0101322 Cited by: §4.3.
- Cometary water expansion velocity from OH line shapes. A&A 467 (2), pp. 729–735. External Links: Document, astro-ph/0702193 Cited by: §4.2.
- Observations of the Hydroxyl Radical in C/2013 US10 (Catalina) at 18 cm Wavelength. AJ 154 (6), pp. 249. External Links: Document Cited by: §1, §4.3.
- Performance of FAST with an Ultra-Wide Bandwidth Receiver at 500-3300 MHz. Research in Astronomy and Astrophysics 23 (7), pp. 075016. External Links: Document, 2304.11895 Cited by: Appendix A, §2, §3.
- Understanding the Activity Performance of Comets 38P/Stephan-Oterma, 64P/Swift-Gehrels and C/2017 M4 (ATLAS) through Broadband Photometric Observations. Research in Astronomy and Astrophysics 24 (8), pp. 085013. External Links: Document Cited by: §1.
Appendix A From Antenna temperature to Flux for FAST UWB data
@@
Here, represents the telescope’s efficiency, and is the theoretical gain based on the geometrical illumination area. We adopt the value of the Gain at the 1650 MHz from (Zhang et al. 2023), according to the zenith angle (ZA) which they detected derive the corresponding 19.6 K Jy-1, then get the actual gain through the relationship between ZA and (Jiang et al. 2019). The result was shown in Table 2.
@@ Appendix B The OH production rate with quenching
When considering the quenching effect, we assume that the inversion factor becomes for , so that OH radicals within this region do not contribute to the observable 18-cm signal (Despois et al. 1981), and the factor can be written as:
Where the is the column density distribution after subtracting the sphere with radius . The quenching radius can be expressed as = (e.g., Li et al. 2025). Adopting following Li et al. (2025) results in unstable convergence. We therefore adopt a reduced value of to ensure convergence. With this treatment, the derived values for October 23, 26 and November 4, 5 are , , , and s-1, respectively.